Can you please put y=3x+6 into a word problem
Answer by jim_thompson5910(35256)


Let's say that you have 6 movies on your shelf. Each week you add on 3 more movies, so that after x weeks, you have added a total of 3x movies (3x = 3*x = 3 times x). In total, you have 3x+6 movies. If y is the number of movies you have after x weeks, then y = 3x+6

If you wanted to figure out how many movies you have say after 3 weeks, then you'd plug in x = 3 to get y = 3*x+6 = 3*3+6 = 9+6 = 15 so you have 15 movies after 3 weeks.

If you wanted to figure out how many weeks it takes to get a certain number of movies, say 21 movies, then plug y = 21 into the equation and solve for x
y = 3x+6
21 = 3x+6
21-6 = 3x+6-6
15 = 3x
3x = 15
3x/3 = 15/3
x = 5
So it takes you 5 weeks to have a total of 21 movies